# AnySR: Realizing Image Super-Resolution as Any-Scale, Any-Resource

### v1

@@ Abstract
In an effort to improve the efficiency and scalability of single-image super-resolution (SISR) applications, we introduce AnySR, to rebuild existing arbitrary-scale SR methods into any-scale, any-resource implementation. As a contrast to off-the-shelf methods that solve SR tasks across various scales with the same computing costs, our AnySR innovates in: 1) building arbitrary-scale tasks as any-resource implementation, reducing resource requirements for smaller scales without additional parameters; 2) enhancing any-scale performance in a feature-interweaving fashion, inserting scale pairs into features at regular intervals and ensuring correct feature/scale processing. The efficacy of our AnySR is fully demonstrated by rebuilding most existing arbitrary-scale SISR methods and validating on five popular SISR test datasets. The results show that our AnySR implements SISR tasks in a computing-more-efficient fashion, and performs on par with existing arbitrary-scale SISR methods. For the first time, we realize SISR tasks as not only any-scale in literature, but also as any-resource. Code is available at https://github.com/CrispyFeSo4/AnySR.
Keywords Super-Resolution  $\cdot$ Any-Scale  $\cdot$ Any-Resource
1 Introduction
Single image super-resolution (SISR) is the process of reconstructing a low-resolution (LR) image into a high-resolution (HR) one rich in detail. The ill-posed nature has made SISR one of the most challenging tasks in low-level computer vision. The past decades have witnessed the advent of many classic studies that continuously address such a challenge. Typical works such as EDSR [1], RCAN [2], and RDN [3], serving as network backbones, effectively extract both texture and semantic information from LR images and lay the foundation for HR reconstruction. The pixel-shuffle upsampling method [4] takes a step further to enhance the image reconstruction quality. Standing upon the shoulders of these pioneering works, the majority of SISR methods [1, 2, 3] adheres to a straightforward and efficient paradigm in which features are extracted by feature extractors, followed by upscaling modules for reconstruction.
@@ 1 Introduction
Albeit the encouraging achievement, most of the above methods are confined to addressing the image reconstruction of a fixed upsampling scale and are therefore stuck in repeatedly rebuilding an SR network when the fully trained model cannot accommodate the resolutions in demand. Actually, there is a significant demand for multi-scale SR tasks. Training individual models for respective resolutions seems to be a great waste of computing resources and fails to adapt to real-time online deployment. Reflecting on this, many recent researchers have shifted their focus to integrating any-scale tasks into only one integrated SR model. For example, Meta-SR [9] learns to predict coordinates and scale-oriented convolutional weights to generate HR images. Taking into consideration the neighborhoods in the LR image and a scale-related cell size, the subsequent LIIF [10] utilizes an MLP to predict the pixel values of the queried coordinates in an HR one. To enhance the representational capabilities, LTE [11] further introduces a local texture estimator that transforms coordinate information into the Fourier domain. The very recent SRNO [12] and OPE-SR [13] enable HR tasks of continuous scales by respectively incorporating Neural Operators [14] and introducing an orthogonal position encoding upscale module.
@@
To this end, we propose AnySR in this paper, a general method with “SR” in the method’s name indicating not only Super-Resolution task but also any-Scale and any-Resource implementation. To be more concrete, we reorganize scale information into multiple sub-groups, each of which consists of similar scale sizes. In this context, we construct multiple networks of various complexities so that in situations where computational resources are constrained, lower-complexity ones are distributed to accomplish SISR tasks of smaller scales for the sake of inference efficiency, while higher-complexity ones are oriented to solving higher-scale SISR tasks on account of better performance. Further, to circumvent the introduction of cumbersome parameters, we train and infer all networks in a parameter-sharing manner [20, 21] in which the weights of smaller networks become parts of larger ones. Such a fashion also conveniences the case of abundant computational resources, where opting for the most complex network produces superior reconstruction results for all scales. Importantly, unlike traditional methods [20, 21] that forward a batch of samples multiple times during training, we forward each batch a time and thus do not enlarge the training costs.
@@
2 Related Work
2.1 Arbitrary-Scale Super-Resolution
@@
2.2 Efficient Image Super-Resolution
In recent years, numerous approaches [42, 43, 44, 45, 46] have emerged to develop efficient super-resolution networks. Building upon the pioneering work of SRCNN [22], FSRCNN [42] substantially accelerates the SISR networks. This acceleration is achieved by taking the original LR image as input without bicubic interpolation, using smaller sizes of convolution kernels, and incorporating a deconvolutional layer at the final stage of the network for upsampling. LapSRN [47] incrementally reconstructs the sub-band residuals of HR images through the utilization of the Laplacian pyramid. CARN [43] achieves improved efficiency by ingeniously designing cascading residual networks with group convolution. IMDN [48] presents information multi-distillation blocks featuring a contrast-aware channel attention layer. In parallel, RFDN [44] optimizes the architecture using a feature distillation mechanism through the proposed residual feature distillation block. In the wake of RFDN, RLFN [49] adds more channels to compensate for discarded feature distillation branches and introduces the residual local feature block, resulting in enhanced inference speed and superior performance with fewer parameters. On the other hand, FMEN [45] introduces the re-parameterization technique, expanding the optimization space during training through re-parameterizable building blocks [50], without incurring additional inference time.
@@ 3.1 Preliminaries
As analyzed in Sec. 1, such an arbitrary-scale pipeline constrains the efficiency and scalability of applying SR models. A less complex network can sometimes adequately handle a smaller reconstruction scale. Consequently, allocating a uniform inference overhead for all scales manifests insufficient scalability to accommodate running platforms of time-varying computational resources, and presents poor efficiency when dealing with smaller reconstruction scales. Contemplating this, in Sec. 3.2, we rebuild current methods as an implementation of “any-resource”, to enable the selection of a smaller network in the SR model based on the abundance of computing resources. Then, we continue proposing “feature-interweaving” in Sec. 3.3 to enhance the performance under our “any-resource” implementation.
3.2 Any-Resource Implementation
@@
Algorithm 1 AnySR Training.
0:  LR image $I_{LR}$, HR image $I_{HR}$, sorted and split scale set $S=S_{1}\cup S_{2}\cup...\cup S_{T}$, a pre-trained arbitrary-scale SR model $F$ with weights $\Theta_{F}$, training step $K$.
1:  for $k=1\rightarrow K$ do
2:     $t\sim[1,2,...,T]$;
3:     $(s^{h},s^{w})\sim S_{t}$;
4:     $t\leftarrow T$ with probability $p$;
5:     Forward $I_{LR}$ and $(s^{h},s^{w})$ to $F$ with weights $\Theta_{F}[1:|\Theta_{F_{t}}|]$;
6:     Obtain reconstructed HR image $I^{\prime}_{HR}$ via Eq. (4);
7:     $F\leftarrow F-\beta\nabla\parallel I_{HR}-I^{\prime}_{HR}\parallel_{1}$.               % Only $\Theta_{F}[1:|\Theta_{F_{t}}|]$ are updated.
8:  end for
9:  return  $F$
@@
$\Theta_{F_{1}}=\Theta_{F}[1:|\Theta_{F_{1}}|],\Theta_{F_{2}}=\Theta_{F}[1:|\Theta_{F_{2}}|],...,\Theta_{F_{T}}=\Theta_{F}[1:|\Theta_{F_{T}}|]=\Theta_{F},$
where $\Theta_{F}[1:|\Theta_{F_{t}}|]$ represents the first $|\Theta_{F_{t}}|$ filters of $\Theta_{F}$, which indeed makes $F_{t}$ a subnet of $F$ as $\Theta_{F_{1}}\subset\Theta_{F_{2}}\subset...\subset\Theta_{F_{T}}\subseteq\Theta_{F}$.
@@ 3.3 Any-Scale Enhancement
Feature-Interweaving. Features $f_{t}\in\mathbb{R}^{H\times W\times(\lfloor C_{in}\cdot w_{t}\rfloor)}$ from $F_{t}$ would first go through a global average pooling (GAP) to get a variant $\tilde{f}_{t}\in\mathbb{R}^{\lfloor C_{\mathrm{in}}\cdot w_{t}\rfloor}$ where the scale information $(s^{h},s^{w})\in S_{t}$ is formally injected.
@@
$f_{t}^{\prime}=f_{t}\odot\mathrm{Sigmoid}\big{(}W_{2}[1:\lfloor C_{\mathrm{in}}\cdot w_{t}\rfloor,:]\cdot\mathrm{ReLU}(W_{1}[:,1:\lfloor C_{\mathrm{in}}\cdot w_{t}\rfloor+2]\cdot\bar{f}_{t})\big{)},$ (5)
@@
Though this naive approach facilitates the interaction between features and scale information, two notable limitations, as we analyze, arise: 1) Insufficient scale information. Compared to a total of $\lfloor C_{\mathrm{in}}\cdot w_{t}\rfloor$ channels in image features, where $\lfloor C_{\mathrm{in}}\cdot w_{t}\rfloor\gg 2$ signifies a limited influence of scale information on the weighted features, the scale information takes up only 2 dimensions. 2) Inappropriate scale processing. For the MLP weights $W_{1}\in\mathbb{R}^{2\cdot C_{\mathrm{in}}\times(C_{\mathrm{in}}+2)}$, the sub-weights $W_{1}[:,\lfloor C_{\mathrm{in}}\cdot w_{t}\rfloor+1:\lfloor C_{\mathrm{in}}\cdot w_{t}\rfloor+2]\in\mathbb{R}^{2\cdot C_{\mathrm{in}}\times 2}$ are used to process the scale information $(s^{h},s^{w})\in S_{t}$. For $(s^{h},s^{w})\in S_{t+1}$, the weights to process scales are $W_{1}[:,\lfloor C_{\mathrm{in}}\cdot w_{t+1}\rfloor+1:\lfloor C_{\mathrm{in}}\cdot w_{t+1}\rfloor+2]\in\mathbb{R}^{2\cdot C_{\mathrm{in}}\times 2}$ and $W_{1}[:,\lfloor C_{\mathrm{in}}\cdot w_{t}\rfloor+1:\lfloor C_{\mathrm{in}}\cdot w_{t}\rfloor+2]\in\mathbb{R}^{2\cdot C_{\mathrm{in}}\times 2}$ are now responsible for processing image features. If off-the-shelf methods [10, 52] are still applied here, as a result, they may weaken the impact of scale information and thus fail to well excavate customized features at different scales. Experimental demonstration has been provided in Table 4.
@@
$\begin{split}\bar{f}_{t}=\Big{[}&\tilde{f}_{t}\big{[}1:\lfloor C_{\mathrm{in}}\cdot 1/{\lambda}\rfloor\big{]},s^{h},s^{w},\quad\tilde{f}_{t}\big{[}\lfloor C_{\mathrm{in}}\cdot 1/{\lambda}\rfloor+1:\lfloor C_{\mathrm{in}}\cdot 2/{\lambda}\rfloor\big{]},s^{h},s^{w},\quad\dots,\\ &\tilde{f}_{t}\big{[}\lfloor C_{\mathrm{in}}\cdot(i-1)/{\lambda}\rfloor+1:\lfloor C_{\mathrm{in}}\cdot{i}/{\lambda}\rfloor\big{]},s^{h},s^{w},\quad\dots,\quad\tilde{f}_{t}\big{[}\lfloor C_{\mathrm{in}}\cdot\lfloor\lambda\cdot w_{t}\rfloor/{\lambda}\rfloor+1:\big{]}\Big{]}.\end{split}$ (6)
Then, weights $W_{1}$ in the two-layer MLP becomes $W_{1}\in\mathbb{R}^{2\cdot C_{\mathrm{in}}\times(C_{\mathrm{in}}+2\cdot\lambda)}$, and the weighted features in (5) becomes:
$f_{t}^{\prime}=f_{t}\odot\mathrm{Sigmoid}\big{(}W_{2}\big{[}1:\lfloor C_{\mathrm{in}}\cdot w_{t}\rfloor,:\big{]}\cdot\mathrm{ReLU}(W_{1}\big{[}:,1:\lfloor C_{\mathrm{in}}\cdot w_{t}\rfloor+2\cdot\lfloor\lambda\cdot w_{t}\rfloor\big{]}\cdot\bar{f}_{t})\big{)}.$ (7)
As a consequence, the weights $W_{1}\big{[}:,\lfloor C_{\mathrm{in}}\cdot{i}/{\lambda}\rfloor+2i-1:\lfloor C_{\mathrm{in}}\cdot{i}/{\lambda}\rfloor+2i\big{]}\in\mathbb{R}^{2\cdot C_{\mathrm{in}}\times 2}$, in which $i=1,2,...,\lfloor\lambda\cdot w_{t}\rfloor$ consistently deal with the scale information whatever the specific value of $t$. In conclusion, our feature-interweaving overcomes the issues of traditional methods [10, 52] by taking into account two key operations: 1) repeating the scale pair $(s^{h},s^{w})$ $\lambda$ times to ensure sufficient scale information for network $F_{t}$; 2) inserting scale pairs into features in regular intervals to ensure the correct feature/scale processing. The efficacy of our feature-interweaving will be shown in Sec. 4.
Table 1: PSNR (dB) of existing arbitrary-scale SR methods and their AnySR variants (through different subnets) highlighted by . The indicates our re-implementation. Following [13], we highlight AnySR performance by blue if less than 0.15db of the vanilla method, and red if better.
@@
RDN MetaSR 38.22 34.63 32.38 33.98 30.54 28.78 32.33 29.26 27.71 32.92 28.82 26.55 39.18 34.14 31.03
@@
4 Experimental Results
4.1 Experiment Settings
Training Details. We primarily rely on existing well-established arbitrary-scale SR models, including Meta-SR [9], LIIF [10], ArbSR [51], and SRNO [12], for AnySR rebuilding. Without loss of generality, experiments on top of two famous feature extraction backbones, including EDSR [1] and RDN [3], are conducted to validate the universality of our AnySR. During specific implementation, we configure the network number $T=4$ and $\{w_{i}\}^{T}_{i=1}=\{0.5,0.7,0.9,1.0\}$, that is, the smallest scale group $S_{1}$ enables a reduction in inference cost by up to 50% compared with the original. The reset probability $p$ in Line 4 of Algorithm 1 is configured at 0.6, and the hyper-parameter $\lambda$ in feature-interweaving is set as 4 for EDSR and 8 for RDN. Accordingly, the scale groups $\mathcal{S}_{1}=\{1.1,1.2,...,1.7\}$, $\mathcal{S}_{2}=\{1.8,1.9,...,2.5\}$, $\mathcal{S}_{3}=\{2.6,2.7,...,3.2\}$ and $\mathcal{S}_{4}=\{3.3,3.4,...,4.0\}$. To ensure a fair performance evaluation, we keep identical settings and configurations used for the original models in our training process. The patch size is $50\times 50$ for Meta-SR and ArbSR, $48\times 48$ for LIIF, and $128\times 128$ for SRNO, with a batch size of 8 per GPU for EDSR and 4 per GPU for RDN. We train AnySR for 500 epochs from the pre-trained models, with an initial learning rate $10^{-5}$ decayed by 0.5 every 200 epochs. We employ an $\ell_{1}$ loss [1] and the Adam optimizer [58].
Table 2: PSNR (dB) of existing arbitrary-scale SR methods and AnySR-retrained version (through the largest network) denoted by . The indicates our re-implementation. Following [13], we highlight AnySR performance by blue if less than 0.15db of the vanilla method, and red if better.
EDSR MetaSR 37.89 34.35 32.08 33.57 30.29 28.52 32.14 29.07 27.54 32.07 28.10 25.97 38.36 33.39 30.37
RDN MetaSR 38.22 34.63 32.38 33.98 30.54 28.78 32.33 29.26 27.71 32.92 28.82 26.55 39.18 34.14 31.03

### v2

@@ Abstract
In an effort to improve the efficiency and scalability of single-image super-resolution (SISR) applications, we introduce AnySR, to rebuild existing arbitrary-scale SR methods into any-scale, any-resource implementation. As a contrast to off-the-shelf methods that solve SR tasks across various scales with the same computing costs, our AnySR innovates in: 1) building arbitrary-scale tasks as any-resource implementation, reducing resource requirements for smaller scales without additional parameters; 2) enhancing any-scale performance in a feature-interweaving fashion, inserting scale pairs into features at regular intervals and ensuring correct feature/scale processing. The efficacy of our AnySR is fully demonstrated by rebuilding most existing arbitrary-scale SISR methods and validating on five popular SISR test datasets. The results show that our AnySR implements SISR tasks in a computing-more-efficient fashion, and performs on par with existing arbitrary-scale SISR methods. For the first time, we realize SISR tasks as not only any-scale in literature, but also as any-resource. Our code is available at https://github.com/CrispyFeSo4/AnySR.
Index Terms:
Super-Resolution Any-Scale Any-Resource
I Introduction
Single image super-resolution (SISR) is the process of reconstructing a low-resolution (LR) image into a high-resolution (HR) one rich in detail. The ill-posed nature has made SISR one of the most challenging tasks in low-level computer vision. The past decades have witnessed the advent of many classic studies that continuously address such a challenge. Typical works such as EDSR [1], RCAN [2], and RDN [3], serving as network backbones, effectively extract both texture and semantic information from LR images and lay the foundation for HR reconstruction. The pixel-shuffle upsampling method [4] takes a step further to enhance the image reconstruction quality. Standing upon the shoulders of these pioneering works, the majority of SISR methods [1, 2, 3] adhere to a straightforward and efficient paradigm in which features are extracted by feature extractors, followed by upscaling modules for reconstruction.
@@ I Introduction
Albeit the encouraging achievement, most of the above methods are confined to addressing the image reconstruction of a fixed upsampling scale, and are therefore stuck in repeatedly rebuilding an SR network, when the fully trained model cannot accommodate the resolutions in demand. Actually, there is a significant demand for multi-scale SR tasks. Training individual models for respective resolutions seems to be a great waste of computing resources and fails to adapt to real-time online deployment. Reflecting on this, many recent researchers have shifted their focus to integrating any-scale tasks into only one integrated SR model. For example, Meta-SR [9] learns to predict coordinates and scale-oriented convolutional weights to generate HR images. Taking into consideration the neighborhoods in the LR image and a scale-related cell size, the subsequent LIIF [10] utilizes an MLP to predict the pixel values of the queried coordinates in an HR one. To enhance the representational capabilities, LTE [11] further introduces a local texture estimator that transforms coordinate information into the Fourier domain. The very recent SRNO [12] and OPE-SR [13] enable HR tasks of continuous scales by respectively incorporating Neural Operators [14] and introducing an orthogonal position encoding upscale module.
@@
To this end, we propose AnySR in this paper, a general method with “SR” in the method’s name indicating not only Super-Resolution task but also any-Scale and any-Resource implementation. To be more concrete, we reorganize scale information into multiple sub-groups, each of which consists of similar scale sizes. In this context, we design multiple networks with varying levels of complexity. For cases where computational resources are limited, networks with lower complexity are assigned to handle smaller-scale SISR tasks to ensure inference efficiency, while more complex networks are used for larger-scale SISR tasks to achieve better performance. Further, to circumvent the introduction of cumbersome parameters, we train and infer all networks in a parameter-sharing manner [20, 21] in which the weights of smaller networks become parts of larger ones. Such a fashion also conveniences the case of abundant computational resources, where opting for the most complex network produces superior reconstruction results for all scales. Importantly, unlike traditional methods [20, 21] that forward a batch of samples multiple times during training, we forward each batch a time and thus do not enlarge the training costs.
@@
II Related Work
II-A Arbitrary-Scale Super-Resolution
@@
II-B Efficient Image Super-Resolution
In recent years, numerous approaches [42, 43, 44, 45, 46] have emerged to develop efficient super-resolution networks. Building upon the pioneering work of SRCNN [22], FSRCNN [42] substantially accelerates the SISR networks. This acceleration is achieved by taking the original LR image as input without bicubic interpolation, using smaller sizes of convolution kernels, and incorporating a deconvolutional layer at the final stage of the network for upsampling. LapSRN [47] incrementally reconstructs the sub-band residuals of HR images through the utilization of the Laplacian pyramid. CARN [43] achieves improved efficiency by ingeniously designing cascading residual networks with group convolution. IMDN [48] presents information multi-distillation blocks featuring a contrast-aware channel attention layer. In parallel, RFDN [44] optimizes the architecture using a feature distillation mechanism through the proposed residual feature distillation block. In the wake of RFDN, RLFN [49] adds more channels to compensate for discarded feature distillation branches and introduces the residual local feature block, resulting in enhanced inference speed and superior performance with fewer parameters. On the other hand, FMEN [45] introduces the re-parameterization technique, expanding the optimization space during training through re-parameterizable building blocks [50], without incurring additional inference time. HAT [38] proposes same-task pre-training, which directly performs pre-training on a larger-scale dataset such as ImageNet [51] based on the same task. Although it requires careful control of the pre-training iteration count and the use of an appropriately small learning rate, this approach is straightforward and leads to significant performance improvements.
@@ III-A Preliminaries
As analyzed in Sec. I, such an arbitrary-scale pipeline constrains the efficiency and scalability of applying SR models. A less complex network can sometimes adequately handle a smaller reconstruction scale. Consequently, allocating a uniform inference overhead for all scales manifests insufficient scalability to accommodate running platforms of time-varying computational resources, and presents poor efficiency when dealing with smaller reconstruction scales. Contemplating this, in Sec. III-B, we rebuild current methods as an implementation of “any-resource”, to enable the selection of a smaller network in the SR model based on the abundance of computing resources. Then, we continue proposing “feature-interweaving” in Sec. III-C to enhance the performance under our “any-resource” implementation.
Algorithm 1 AnySR Training.
0:  LR image $I_{\mathrm{LR}}$, HR image $I_{\mathrm{HR}}$, sorted and split scale set $S=S_{1}\cup S_{2}\cup...\cup S_{T}$, a pre-trained arbitrary-scale SR model $F$ with weights $\Theta_{F}$, training step $K$.
1:  for $k=1\rightarrow K$ do
2:     $t\sim[1,2,...,T]$;
3:     $(s^{h},s^{w})\sim S_{t}$;
4:     $t\leftarrow T$ with probability $p$;
5:     Forward $I_{\mathrm{LR}}$ and $(s^{h},s^{w})$ to $F$ with weights $\Theta_{F}[1:|\Theta_{F_{t}}|]$;
6:     Obtain reconstructed HR image $I^{\prime}_{\mathrm{HR}}$ via Eq. (5);
7:     $F\leftarrow F-\beta\nabla\parallel I_{\mathrm{HR}}-I^{\prime}_{\mathrm{HR}}\parallel_{1}$.   % $\Theta_{F}[1:|\Theta_{F_{t}}|]$ are updated.
8:  end for
9:  return  $F$
III-B Any-Resource Implementation
@@
$\Theta_{F_{1}}$ $=\Theta_{F}[1:|\Theta_{F_{1}}|],$ (4)
$\Theta_{F_{2}}$ $=\Theta_{F}[1:|\Theta_{F_{2}}|],$
$\vdots$
$\Theta_{F_{T}}$ $=\Theta_{F}[1:|\Theta_{F_{T}}|]=\Theta_{F},$
where $\Theta_{F}[1:|\Theta_{F_{t}}|]$ represents the first $|\Theta_{F_{t}}|$ filters of $\Theta_{F}$, making $F_{t}$ a subnet of $F$ as $\Theta_{F_{1}}\subset\Theta_{F_{2}}\subset...\subset\Theta_{F_{T}}\subseteq\Theta_{F}$.
@@ III-C Any-Scale Enhancement
Feature-Interweaving.
Features $f_{t}\in\mathbb{R}^{H\times W\times(\lfloor C_{in}\cdot w_{t}\rfloor)}$ from $F_{t}$ would first go through a global average pooling (GAP) to get a variant $\tilde{f}_{t}\in\mathbb{R}^{\lfloor C_{\mathrm{in}}\cdot w_{t}\rfloor}$ where the scale information $(s^{h},s^{w})\in S_{t}$ is formally injected.
@@
$f_{t}^{\prime}$ $=f_{t}\odot\mathrm{Sigmoid}\big{(}W_{2}[1:\lfloor C_{\mathrm{in}}\cdot w_{t}\rfloor,:]\cdot$ (6)
$\quad\mathrm{ReLU}(W_{1}[:,1:\lfloor C_{\mathrm{in}}\cdot w_{t}\rfloor+2]\cdot\bar{f}_{t})\big{)},$
@@
Though this naive approach facilitates the interaction between features and scale information, two notable limitations, as we analyze, arise: 1) Insufficient scale information. Compared to a total of $\lfloor C_{\mathrm{in}}\cdot w_{t}\rfloor$ channels in image features, where $\lfloor C_{\mathrm{in}}\cdot w_{t}\rfloor\gg 2$ signifies a limited influence of scale information on the weighted features, the scale information takes up only 2 dimensions. 2) Inappropriate scale processing. For the MLP weights $W_{1}\in\mathbb{R}^{2\cdot C_{\mathrm{in}}\times(C_{\mathrm{in}}+2)}$, the sub-weights $W_{1}[:,\lfloor C_{\mathrm{in}}\cdot w_{t}\rfloor+1:\lfloor C_{\mathrm{in}}\cdot w_{t}\rfloor+2]\in\mathbb{R}^{2\cdot C_{\mathrm{in}}\times 2}$ are used to process the scale information $(s^{h},s^{w})\in S_{t}$. For $(s^{h},s^{w})\in S_{t+1}$, the weights to process scales are $W_{1}[:,\lfloor C_{\mathrm{in}}\cdot w_{t+1}\rfloor+1:\lfloor C_{\mathrm{in}}\cdot w_{t+1}\rfloor+2]\in\mathbb{R}^{2\cdot C_{\mathrm{in}}\times 2}$ and $W_{1}[:,\lfloor C_{\mathrm{in}}\cdot w_{t}\rfloor+1:\lfloor C_{\mathrm{in}}\cdot w_{t}\rfloor+2]\in\mathbb{R}^{2\cdot C_{\mathrm{in}}\times 2}$ are now responsible for processing image features. If off-the-shelf methods [10, 53] are still applied here, as a result, they may weaken the impact of scale information and thus fail to well excavate customized features at different scales. Experimental demonstration has been provided in Table VII.
@@
$\begin{split}\bar{f}_{t}=\Big{[}&\tilde{f}_{t}\big{[}1:\lfloor C_{\mathrm{in}}\cdot 1/{\lambda}\rfloor\big{]},s^{h},s^{w},\\ &\tilde{f}_{t}\big{[}\lfloor C_{\mathrm{in}}\cdot 1/{\lambda}\rfloor+1:\lfloor C_{\mathrm{in}}\cdot 2/{\lambda}\rfloor\big{]},s^{h},s^{w},\\ &\dots,\\ &\tilde{f}_{t}\big{[}\lfloor C_{\mathrm{in}}\cdot(i-1)/{\lambda}\rfloor+1:\lfloor C_{\mathrm{in}}\cdot{i}/{\lambda}\rfloor\big{]},s^{h},s^{w},\dots,\\ &\tilde{f}_{t}\big{[}\lfloor C_{\mathrm{in}}\cdot\lfloor\lambda\cdot w_{t}\rfloor/{\lambda}\rfloor+1:\big{]},s^{h},s^{w}\Big{]}.\end{split}$ (7)
Notably, the scale pair $(s^{h},s^{w})$ is not always appended at the end of $\bar{f}_{t}$. For Convenience, we define the fixed interval between these insertions as $C_{fix}=C_{in}/\lambda$, and refer to the last portion of the feature $\bar{f}_{t}$ as $C_{last}$. When $C_{last}<C_{fix}$, the scale pair is not appended, ensuring correct feature and scale processing. Conversely, when $C_{last}=C_{fix}$, the scale pair is appended to the end of the feature.
Then, weights $W_{1}$ in the two-layer MLP become $W_{1}\in\mathbb{R}^{2\cdot C_{\mathrm{in}}\times(C_{\mathrm{in}}+2\cdot\lambda)}$, and the weighted features in Eq. (6) is:
$\begin{split}f_{t}^{\prime}=&f_{t}\odot\mathrm{Sigmoid}\big{(}W_{2}\big{[}1:\lfloor C_{\mathrm{in}}\cdot w_{t}\rfloor,:\big{]}\cdot\\ &\mathrm{ReLU}(W_{1}\big{[}:,1:\lfloor C_{\mathrm{in}}\cdot w_{t}\rfloor+2\cdot\lfloor\lambda\cdot w_{t}\rfloor\big{]}\cdot\bar{f}_{t})\big{)}.\end{split}$ (8)
As a consequence, the weights $W_{1}\big{[}:,\lfloor C_{\mathrm{in}}\cdot{i}/{\lambda}\rfloor+2i-1:\lfloor C_{\mathrm{in}}\cdot{i}/{\lambda}\rfloor+2i\big{]}\in\mathbb{R}^{2\cdot C_{\mathrm{in}}\times 2}$, in which $i=1,2,...,\lfloor\lambda\cdot w_{t}\rfloor$ consistently deal with the scale information whatever the specific value of $t$. In conclusion, our feature-interweaving overcomes the issues of traditional methods [10, 53] by taking into account two key operations: 1) repeating the scale pair $(s^{h},s^{w})$ $\lambda$ times to ensure sufficient scale information for network $F_{t}$; 2) inserting scale pairs into features in regular intervals to ensure the correct feature/scale processing. The efficacy of our feature-interweaving will be shown in Sec. IV.
TABLE I: PSNR (dB) and LPIPS comparison of existing arbitrary-scale SR methods and their AnySR variants, highlighted by (through different subnets) and (AnySR-retrained version via the largest network). denotes our re-implementation. Following [13], AnySR performance is highlighted in blue if within 0.15dB (PSNR) / 0.01 (LPIPS) of the vanilla method, and red if improved.
@@
[…]
MetaSR 0.022 0.054 0.078 0.040 0.089 0.126 0.055 0.102 0.139 0.044 0.100 0.147 0.012 0.036 0.062
MetaSR 0.023 0.053 0.076 0.041 0.087 0.125 0.055 0.101 0.137 0.044 0.098 0.144 0.012 0.034 0.059
MetaSR 0.022 0.053 0.076 0.040 0.087 0.125 0.055 0.101 0.137 0.042 0.096 0.144 0.012 0.034 0.059
LIIF 0.040 0.093 0.135 0.071 0.164 0.231 0.113 0.231 0.310 0.049 0.124 0.187 0.074 0.104 0.136
LIIF 0.041 0.098 0.143 0.074 0.169 0.238 0.113 0.237 0.320 0.053 0.132 0.199 0.075 0.106 0.141
LIIF 0.041 0.096 0.139 0.072 0.167 0.235 0.113 0.235 0.316 0.049 0.125 0.189 0.075 0.105 0.137
ArbSR 0.022 0.052 0.075 0.038 0.085 0.123 0.054 0.097 0.136 0.038 0.088 0.138 0.011 0.033 0.059
ArbSR 0.023 0.052 0.074 0.041 0.085 0.123 0.055 0.097 0.135 0.042 0.090 0.139 0.011 0.033 0.059
ArbSR 0.022 0.052 0.074 0.038 0.085 0.123 0.054 0.097 0.135 0.037 0.088 0.139 0.011 0.033 0.059
SRNO 0.040 0.092 0.137 0.070 0.162 0.226 0.111 0.229 0.303 0.045 0.118 0.175 0.015 0.045 0.077
SRNO 0.041 0.092 0.138 0.072 0.164 0.227 0.112 0.230 0.304 0.048 0.120 0.176 0.016 0.046 0.078
SRNO 0.040 0.093 0.138 0.071 0.163 0.227 0.112 0.230 0.304 0.047 0.119 0.176 0.016 0.046 0.078
[…]
MetaSR 38.22 34.63 32.38 33.98 30.54 28.78 32.33 29.26 27.71 32.92 28.82 26.55 39.18 34.14 31.03